After Aidan validated that model inputs were creating physical parameter changes, we proceeded with some last few checks before closing up. Notes:
- Aidan set up a helpful script turning laser power on and off, and strip tool to follow PD and monitor PD signals. A strip tool chart was used to confirm that there was no loss of functionality or alignment during pump down.
- Checked Heater function and Workpiece RTD response - all good.
- Confirmed alignment by steering at periscope output mirror and watching PD voltage.
At this point Aidan gave a green light for pumpdown prep, and to start pumpdown and start cooldown. Notes:
- Stephen disconnected the PD connector accidentally while trying to add strain relief (the irony is palpable). Reattachment of connector didn't seem to affect any signal levels.
- During installation of radiation shield lids, shields became misaligned and PD signal fell (presumably due to clipping). Recovered previous signal levels by realligning outer shield.
- Double checked that everything seemed good to go, no issues!
- Timeline of pumpdown and cooldown:
17:32 - Pumpdown started.
17:47 - Turbo spin up started (pump station delay parameter).
17:55 - Pressure dropped below 1 mTorr, so I started cryocooler.
18:05 - Healthy so far - Pressure had come down half a decade more, and Coldhead RTD was reading 235 K.